Operations report to EB

  • Data-taking
    • The current user cycle (2015/04) has seen
      • alignment and calibration data-taking
      • return and comissioning of the DS PSU
      • data-taking with vacuum absorber (no magnetic fields)
      • installation of the LiH absorber (in 2 days!)
      • data-taking with LiH absorber (no magnetic fields) due to start today/tomorrow
      • we've negotiated an increase in ISIS beamloss trip levels to 2.5 This should allow
        us to run at 2.0 rather than 1.3 (which is what we've been averaging) and help with the data rate.
    • We are running 16 hours / 7 days a week during this user cycle
    • All shifts have been allocated.
  • DC status
    • DCs have all completed STFC safety training and been confirmed in their positions by head of PPD
    • DCs in Feb have been Paul Hodgson, Victoria and Craig.
    • Co-ordination and communication between DCs is good
    • Formal DC/MOM handover procedure is defined and works well.
  • MOM status
  • Staffing issues
    • Small size of the BLOC pool almost affected data-taking this user cycle
    • We've (re)trained some new BLOCs (Jan, Adam and Ken) which will help
    • BLOCs have been assigned some shift credit as an incentive
    • Pool is still, in my opinion, too small.
  • Next cycle : (April/May - 2016/01)
    • Next user cycle will be mostly given to magnet and power system work.
    • We may take data on the weekends in April to bolster the zero-absorber data-set.
  • Cycle 2016/02 : (June/July)
    • What we do here very much depends on what is ready
    • The current plan is to take energy loss data with LiH (or vacuum) and magnetic fields (SSU (E-C-E) + FC + SSD (E-C-E))
    • We need to understand if (and when) we power M2 in SSD for a cooling measurement

Updated by Boyd, Steven about 8 years ago ยท 9 revisions